# ClassInvGen: Class Invariant Synthesis using Large Language Models

### v1

@@ Abstract.
In this work, we describe ClassInvGen, a method for co-generating executable class invariants and test inputs to produce high-quality class invariants for a mainstream language such as C++, leveraging LLMs’ ability to synthesize pure functions. We show that ClassInvGen outperforms a pure LLM-based technique to generate specifications (from code) as well as prior data-driven invariant inference techniques such as Daikon. We contribute a benchmark of standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated specifications using tests and mutants. We also demonstrate its applicability to real-world code by performing a case study on several classes within a widely used and high-integrity C++ codebase.
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
@@
These program invariants help make explicit the assumptions on the rest of the code helping modular review, reasoning, and analysis. Program invariants are useful for several aspects of software construction and maintenance during the lifetime of a program. First, executable program invariants can be enforced at runtime, thus serving as an early indicator of state corruption helping root causing and stopping a program instead of producing unexpected values. Runtime invariants serve as additional test oracles to amplify testing efforts to catch subtle bugs related to state corruption; this in turn helps with regression testing as the program evolves to satisfy new requirements. The utility of program invariants has led to design-by-contract in languages such as Eiffel (Meyer, 1992), as well as support in other languages such as Java (JML (Burdy et al., 2005)) and .NET (Code Contracts (Fähndrich, 2010)). Further, for languages that support static formal verification (e.g., Dafny (Leino, 2010), Verus (Lattuada et al., 2023), F* (Swamy et al., 2011), Frama-C (Kirchner et al., 2015)), these invariants can serve as the specification as well as help make formal verification scale by making verification modular and scalable. Unfortunately, invariants are underutilized because they require additional work and are sometimes difficult to write, so it would be useful to find a way to generate them automatically.
@@
  • Class invariants are crucial to maintain the integrity of data structures and help point to state corruption that may manifest much later within the class or in the clients. Documenting such implicit contracts can greatly aid the understanding for maintainers of the class.
  • Class invariants often form important parts of preconditions and postconditions for high-integrity data structures. Capturing such invariants in a method and asserting them in preconditions and postconditions help reduce bloat in the specifications.
  • Class invariants are challenging for users to write as writing them requires global reasoning across all the public methods for the class.
For example, consider the class for a doubly linked list as implemented in Z3 Satisfiability Modulo Theories (SMT) code [^1] https://github.com/Z3Prover/z3/blob/master/src/util/dlist.h.
@@ 14 SASSERT(other->invariant());
Figure 1. An invariant in the doubly linked list class in Z3.
@@
Synthesizing program invariants has been an active line of research, with both static and dynamic analysis-based approaches. Static analysis approaches based on variants of abstract interpretation (Cousot and Cousot, 1977) and interpolation (Henzinger et al., 2004) create invariants that are sound by construction. However, such techniques do not readily apply to mainstream programming languages with complex language constructs or require highly specialized methods that do not scale to large modules, since the invariants need to be additionally provably inductive to be retained. On the other hand, Daikon (Ernst et al., 2007) and successors learn invariants dynamically by instantiating a set of templates and retaining the predicates that hold on concrete test cases. While applicable to any mainstream language, it is well known that Daikon generated invariants overfit the test cases and are not sound for all test cases (Polikarpova et al., 2009). Recent works have studied fine-tuning large language models (LLMs) to learn program invariants (Pei et al., 2023) but these methods inherit the same limitation from Daikon because their training data consists of Daikon generated invariants. More importantly, the approach has not been evaluated on stateful classes to construct class invariants.
Recent work on prompting LLMs such as GPT-4 to generate program invariants for mainstream languages (Endres et al., 2024; Greiner et al., 2024; Ma et al., 2024) has been used to generate preconditions, postconditions, and loop invariants, but these methods do not readily extend to generating class invariants. Further, these methods cannot construct expressive invariants that require iterating over complex linked data structures (such as in Figure 1) other than simple arrays.
@@
We leverage the framework proposed by Endres et al. (Endres et al., 2024) to evaluate the test-set correctness and completeness given a set of hidden validation tests and mutants. We contribute a new benchmark comprising standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated invariants (Section 4.1). We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating program invariants from code (Sections 4.34.4) as well as prior data-driven invariant inference techniques such as Daikon (Section 4.5). We also demonstrate its applicability for real-world code by performing a case study on ema, dlist, heap, hashtable, permutation, scoped_vector, and the most complex class, bdd_manager within the Z3 SMT solver codebase; the developers of the codebase confirmed most of the new invariants proposed by ClassInvGen for these modules (Section 5).
@@
  • Present new techniques of invariant-test co-generation by combining simple static analysis with LLM and implement an end-to-end prototype (Section 3).
  • Introduce a high-quality ClassInvGen-instrumented benchmark for evaluating object invariants (Section 4.1).
  • Investigate LLM-assisted class invariant synthesis (Sections 4.34.5).
  • Conduct a case study on Z3 bdd_manager module using ClassInvGen (Section 5).
2. Running Example: AVL Tree
Throughout this paper, AvlTree (Wikipedia contributors, 2024; Dimitrios, 2023) from our benchmark (Section 4.1) will be used as a running example to illustrate the workflow. An AVL tree is a self-balancing binary search tree (BST) where the difference in heights between the left and right subtrees of any node (called the balance factor) is at most 1. This ensures that the tree remains approximately balanced. In this implementation shown in Figure 2, the AvlTree class contains several public methods: insert, remove, contains, clear, height, size, empty, in_order_traversal, pre_order_traversal, post_order_traversal. AvlTree maintains several class invariants determined by the authors of this paper:
@@ 20 T data;
Figure 2. Header file of AvlTree.
3. Approach
Figure 3. Overview of ClassInvGen.
An overview of the ClassInvGen framework is shown in Figure 3. It outlines an automated pipeline for inferring class invariants from source code. ClassInvGen takes a complete source program as input and outputs invariant candidates it has identified with high confidence (called filtered invariants). ClassInvGen starts with a preprocessing step which performs static analysis on the program (Section 3.1.1). Next, an LLM is used to generate candidate invariants and filtering test suites (Section 3.1.2). Then, the code is instrumented to facilitate checking candidate invariants (Section 3.1.3). Finally, ClassInvGen uses generated tests to prune invariant candidates (Section 3.2), and a refinement loop is used to iteratively improve the results (Section 3.3).
3.1. Generation
3.1.1. Preprocessing
As illustrated in Figure 3, the generation phase begins with a static analysis of the source program. ClassInvGen uses a Tree-Sitter-based parser for program preprocessing; Tree-Sitter (Brunsfeld et al., 2024) is a parser generator tool that constructs a syntax tree from source files.
1
2
3
4Function GenerateInvariant(target_class):
5       if target_class.id $\in$ invariants_dict then
6             return invariants_dict[target_class.id]
7
8
9      dep_classes $\leftarrow$ getClassRecursively(target_class)
11
12      foreach class $\in$ rev_topsorted_classes do
13
14            class_code $\leftarrow$ getCodeForClass(class, invariants_dict)
15
16            foreach dep $\in$ getClassDependencies(class) do
17                   // Invariant: dep has been generated invariants for
18                   dep_code $\leftarrow$ getCodeForClass(dep, invariants_dict)
19                   class_code $\leftarrow$ class_code + dep_code
20
21
23
24
25      final_invariants $\leftarrow$ invariants_dict[target_class.id]
26       return final_invariants
27
28
29Function getCodeForClass(class, invariants_dict, include_method_bodies=False):
30       class_text $\leftarrow$ class.get_declaration_text()
31        // Header code
33             class_text $\leftarrow$ invariants_dict[class.id] + class_text
34              // Get generated invariants
35
36
37      if include_method_bodies then
38             // Add methods bodies to class_text if context allows
39
40      return class_text
41
Algorithm 1 Function to generate invariants for target class AST parsed by Tree-Sitter
3.1.2. Generation by LLM
4
6- An assertion in the programming language
9
10Task Description:
13
Figure 4. ClassInvGen Generation system prompt: instruction and task description.
After building the source program AST, ClassInvGen uses LLMs to analyze the class module and infers both invariants for the target class and tests that exercise the class’s implementation as thoroughly as possible. ClassInvGen uses a fixed system prompt that defines class invariants and outlines two main tasks: (1) generating class invariants from the source code, and (2) creating a test suite of valid API calls without specifying expected outputs (Figure 4 and Figure 23 in Appendix). Next, ClassInvGen instantiates a user prompt template (Figure 24 in Appendix) with the actual target class.
From the source program AST, ClassInvGen identifies program dependencies and populates the prompt template with the leaf struct/class. Starting from the leaf nodes, ClassInvGen leverages previously generated invariants by including them in the prompt for later classes. To accommodate the LLM’s context window limit, only the relevant child classes of the current target class are included in the prompt, with method implementations and private fields/methods hidden when necessary. An algorithm for this process is presented in Algorithm 1.
Algorithm 1 presents the invariant generation process for a source program AST. The main function GenerateInvariant takes a target_class and leverages a caching mechanism through invariants_dict to avoid redundant computations (Line 1). The algorithm first collects dependent classes via getClassRecursively (Line 1) and sorts them using reverseToplogicalSort to ensure dependency-aware processing (Line 1). For each class in the sorted order, it constructs the necessary context by obtaining the class code through getCodeForClass (Line 1). For each dependency dep of the current class, it retrieves the dep_code and concatenates it with class_code (Lines 11). The algorithm then generates invariants using generateInvariantWithLLM and stores them in invariants_dict (Line 1). The helper function getCodeForClass constructs class representations by combining the declaration text with any existing invariants from invariants_dict (Lines 11). It optionally includes method bodies based on context window constraints. This approach ensures efficient invariant generation while maintaining all necessary context and dependencies (Line 1). The algorithm concludes by returning final_invariants for the target class, effectively managing the invariant generation process while respecting LLM context limitations.
@@ 3.1.2. Generation by LLM
For the AvlTree example, we begin by instantiating the prompt with Node for annotation, followed by AvlTree, since Node is a subclass of AvlTree, as illustrated in Figure 2. In this specific case, however, Algorithm 1 does not make a difference due to the small size of the source program; the entire AvlTree code fits within the LLM’s context window easily. In contrast, when working with the bdd_manager class in Z3 (around 1700 lines of code), ClassInvGen begins generation with bdd, a subclass of bdd_manager. Algorithm 1 enables ClassInvGen to partition bdd_manager class and outputs meaningful class invariants (see Section 5).
3.1.3. Instrumentation
1bool AvlTree::empty() {
2 check_invariant();
3 auto ret = empty_original();
4 check_invariant();
5 return ret;
6}
7
8bool AvlTree::empty_original() { return n == 0; }
Figure 5. AvlTree Instrumented with invariants
1void AvlTree::check_invariant() {
2 std::function<bool(const std::unique_ptr<Node>&)> is_balanced = [&](const std::unique_ptr<Node>& node) -> bool {
3 if (!node) return true;
4 int left_height = height(node->left);
5 int right_height = height(node->right);
6 if (std::abs(left_height - right_height) > 1) return false;
7 return is_balanced(node->left) && is_balanced(node->right);
8 };
9 assert(is_balanced(root));
10
11}
Figure 6. Example of a correct AvlTree class invariant
1void AvlTree::check_invariant() {
2 assert(height(root) == get_height(root));
3}
Figure 7. Example of an incorrect invariant of AvlTree because there is no get_height method
To check candidate invariants, each public method is automatically instrumented with a check_invariant call at both the start and end of its implementation, as shown in Figure 5. ClassInvGen instruments these invariants to use the compiler to rule out syntactic/type-based errors in the invariants. Each invariant is implemented as a method call to prevent conflicts with local variables within the two invocations to check_invariant. To check a specific invariant, the code for the invariant is plugged in as the implementation for the check_invariant function, and assertions are added to check that the invariant holds, as depicted in Figures 6 and 7. This setup ensures that during pruning, described below, whenever a public API call is made, each invariant candidate is automatically checked.
3.2. Heuristic Pruning
@@
As discussed above, the LLM generates test suites for candidates. ClassInvGen wishes to select the suite that is most effective at identifying inaccurate invariants. To select the most suitable test suite, we employ line coverage as a metric. Line coverage is used to select the best test suite because it serves as a straightforward proxy for test suite completeness. The tests with the highest coverage are considered filtering tests and are used to prune the candidate invariants.
Notably, when generating tests, ClassInvGen focuses on creating valid sequences of API calls to exercise the class under test but does not assert the expected behavior. This approach simplifies the process by eliminating the need for a test oracle, which is unnecessary in the ClassInvGen framework since the purpose of test generation is to filter invariant candidates, not to test the source program directly.
@@
[…]
These filtering tests, which are assumed to contain valid sequences, are then used to identify high-quality invariant candidates, referred to as filtered invariants. If an invariant candidate successfully compiles and runs with the filtering tests, it is designated a filtered invariant and the final output of ClassInvGen.
Figure 6 illustrates a correct filtered invariant, stating that the height of an AvlTree is balanced. In contrast, Figure 7 shows an incorrect invariant that was filtered out by the filtering tests during compilation because it references a get_height method, which is not accessible to the user.
To filter invariants, ClassInvGen compiles and runs candidate invariants with the filtering tests, retaining those that do not cause crashes and adding them to the set of filtered invariants.
1int main() {
2 // Test Case 1
3 // Basic insertions and traversals
4 {
5 AvlTree tree;
6 tree.insert(10);
7 tree.insert(20);
8 tree.insert(5);
9 tree.in_order_traversal();
10 tree.pre_order_traversal();
11 tree.post_order_traversal();
12 }
13 // Test Case 2
14 // Checking size, height, and empty status
15 {
16 AvlTree tree;
17 tree.insert(10);
18 tree.insert(20);
19 tree.insert(5);
20 tree.size();
21 tree.height();
22 tree.empty();
23 }
24 ... // the rest tests are skipped
25}
Figure 8. A test suite generated for AvlTree
3.3. Refinement
Invariants that fail at either compilation or runtime are candidates for refinement loops. For such failing invariants, ClassInvGen implements a feedback loop by collecting compiler output, error messages, and the results of filtering tests, then feeding this information back to the LLM using a dedicated prompt template (Figures 2526 in Appendix) to generate refined invariants. The system prompt for refinement is fixed, while the user prompt is instantiated with the failing invariant, compiler error messages, the target class’s source code, and the filtering tests that caused the failure.
For instance, the GCC error output shown in Figure 9 for the AvlTree candidate invariant in Figure 10 leads to the refined candidate invariant in Figure 11. We set the default threshold for refinement loops to 3, balancing the cost of calling LLMs with the benefit of repairing a single invariant; empirically, this number works well, though it may require adjustment for different codebases. If the invariant still fails after 3 refinement attempts, ClassInvGen discards it and moves on. Invariants are accepted as filtered only after successfully passing all filtered tests.
1avl_tree.cpp: In lambda function:
2avl_tree.cpp:21:16: error: use of ’is_balanced’ before deduction of ’auto’
Figure 9. gcc compiler error messages
1void AvlTree::check_invariant() {
2  auto is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
3 if (!node) return true;
4 if (node->data <= min $|$ node->data >= max) return false;
5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
6 };
7
9}
Figure 10. BST property before refinement: the word auto is causing the compiler to complain
1void AvlTree::check_invariant() {
2  std::function<bool(const std::unique_ptr<Node>&)> is_bst = [&](const std::unique_ptr<Node>& node, const T& min, const T& max) -> bool {
3 if (!node) return true;
4 if (node->data <= min $|$ node->data >= max) return false;
5 return is_bst(node->left, min, node->data) && is_bst(node->right, node->data, max);
6 };
7
9}
Figure 11. BST property after refinement: auto is changed to explicit declarations. This change resolved type ambiguity, allowing the code to compile successfully.
4. Results
4.1. Benchmark
Table 1. Characteristics of the benchmark data structures. # LoC represents lines of code, # methods indicates the number of implemented methods, and # dep. shows the number of dependent classes for each data structure.

### v2

@@ Abstract
In this work, we describe ClassInvGen, a method for co-generating executable class invariants and test inputs to produce high-quality class invariants for a mainstream language such as C++, leveraging LLMs’ ability to synthesize pure functions. We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating specifications (from code) as well as prior data-driven invariant inference techniques such as Daikon. We contribute a benchmark of standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated specifications using tests and mutants. We also demonstrate its applicability to real-world code by performing a case study on several classes within a widely used and high-integrity C++ codebase.
Keywords:
Program Synthesis Large Language Models Class Invariants Formal Verification
1 Introduction
@@
These program invariants help make explicit the assumptions on the rest of the code, helping modular review, reasoning, and analysis. Program invariants are useful for several aspects of software construction and maintenance during the lifetime of a program. First, executable program invariants can be enforced at runtime, where they provide an early indicator of state corruption, help with root causing, and allow a program to halt with an error instead of producing unexpected values. Runtime invariants serve as additional test oracles to amplify testing efforts to catch subtle bugs related to state corruption; this, in turn, helps with regression testing as the program evolves to satisfy new requirements. The utility of program invariants has led to design-by-contract in languages such as Eiffel [32], as well as support in other languages such as Java (JML [5]) and .NET (Code Contracts [12]). Furthermore, for languages that support static formal verification (e.g., Dafny [26], Verus [24], F* [41], Frama-C [20]), invariants can serve as a part of the specification, helping make formal verification modular and scalable. Unfortunately, invariants are underutilized because they require additional work and are sometimes difficult to write, so it would be useful to find a way to generate them automatically.
@@
  • Class invariants are crucial for maintaining the integrity of data structures and help point to state corruption that may manifest much later within the class or in the clients. Documenting such implicit contracts can greatly aid the understanding for maintainers of the class.
  • Class invariants often form important parts of preconditions and postconditions for high-integrity data structures. Encapsulating such invariants and asserting them in preconditions and postconditions helps reduce bloat in the specifications.
  • Class invariants are challenging for users to write, as writing them requires global reasoning across all the public methods for the class.
For example, consider the class in Figure 1 for a doubly-linked list, as implemented in the Z3 SMT solver. [^1] https://github.com/Z3Prover/z3/blob/master/src/util/dlist.h
@@ 14 SASSERT(other->invariant());
Figure 1: An invariant in the doubly linked list class in Z3.
@@
Synthesizing program invariants has been an active line of research, with both static and dynamic analysis-based approaches. Static analysis approaches based on variants of abstract interpretation [7] and interpolation [18] create invariants that are sound by construction. However, such techniques do not readily apply to mainstream programming languages with complex language constructs or require highly specialized methods that do not scale to large modules, since the invariants need to be additionally provably inductive to be retained. On the other hand, Daikon [11] and successors learn invariants dynamically by instantiating a set of templates and retaining the predicates that hold on concrete test cases. While applicable to any mainstream language, it is well known that Daikon-generated invariants overfit the test cases and are not sound for all test cases [39]. Recent works have studied fine-tuning large language models (LLMs), to learn program invariants [38] but these methods inherit the limitations of Daikon because their training data consists of Daikon-generated invariants. More importantly, the approach has not been evaluated on stateful classes to construct class invariants.
Recent work on prompting LLMs such as GPT-4 to generate program invariants for mainstream languages [10, 16, 30] has been used to generate preconditions, postconditions, and loop invariants, but these methods do not readily extend to generating class invariants. These pipelines work at single-loop or single-method scope and validate only scalar, intraprocedural predicates, so they lack the heap-aware, cross-method perspective required to state class invariants. Further, these methods cannot construct expressive invariants that require iterating over complex data structures (such as in Figure 1) other than simple arrays.
@@
We leverage the framework proposed by Endres et al. [10] to evaluate the test-set correctness and completeness given a set of hidden validation tests and mutants. We contribute a new benchmark comprising standard C++ data structures along with a harness that can help measure both the correctness and completeness of generated invariants (Section 4.1). We demonstrate that ClassInvGen outperforms a pure LLM-based technique for generating program invariants from code (Sections 4.34.4) as well as prior data-driven invariant inference techniques such as Daikon (Section 4.5). We also demonstrate its applicability for real-world code by performing a case study on a set of classes in the Z3 SMT solver codebase, including the relatively complex bdd_manager class; the developers of the codebase confirmed most of the new invariants proposed by ClassInvGen for these modules (Section 5).
@@
  • We introduce a new technique for invariant-test co-generation by combining simple static analysis with LLMs and implement an end-to-end prototype (Section 3).
  • We introduce a high-quality ClassInvGen-instrumented benchmark for evaluating object invariants (Section 4.1).
  • We investigate LLM-assisted class invariant synthesis (Sections 4.34.5).
  • We conduct a case study on Z3 class modules using ClassInvGen (Section 5).
ClassInvGen is conceived as a specification-drafting aid: it produces candidate invariants that a developer can accept, refine, or discard, thereby following the long-advocated “human-in-the-loop” paradigm in specification mining [35]. Our aim is not to replace expert judgement, but to accelerate it.
2 Running Example: AVL Tree
Throughout this paper, AvlTree [43, 9] from our benchmark (Section 4.1) will be used as a running example to illustrate the workflow. An AVL tree is a self-balancing binary search tree (BST) where the difference in heights between the left and right subtrees of any node (called the balance factor) is at most 1. This ensures that the tree remains approximately balanced. In this implementation shown in Figure 3 in Section 3.1, the AvlTree class contains several public methods: insert, remove, contains, clear, height, size, empty, in_order_traversal, pre_order_traversal, post_order_traversal. AvlTree maintains several class invariants determined by the authors of this paper:
@@ 2 Running Example: AVL Tree
3 Approach
Figure 2: Overview of ClassInvGen.
An overview of the ClassInvGen framework is shown in Figure 2. It outlines an automated pipeline for inferring class invariants from source code. ClassInvGen takes a complete source program as input and outputs invariant candidates it has identified with high confidence (called filtered invariants). ClassInvGen starts with a preprocessing step which performs static analysis on the program (Section 3.1). Next, an LLM is used to generate candidate invariants and filtering test suites (Section 3.1). Then, the code is instrumented to facilitate checking candidate invariants (Section 3.1). Finally, ClassInvGen uses generated tests to prune invariant candidates (Section 3.2), and a refinement loop is used to iteratively improve the results (Section 3.3).
3.1 Generation
3.1.1 Preprocessing
As illustrated in Figure 2, the generation phase begins with a static analysis of the source program. ClassInvGen uses a Tree-Sitter-based parser for program preprocessing; Tree-Sitter [3] is a parser generator tool that constructs a syntax tree from source files.
1
2
3
4Function GenerateInvariant(target_class):
5       if target_class.id $\in$ invariants_dict then
6             return invariants_dict[target_class.id]
7
8
9      dep_classes $\leftarrow$ getClassRecursively(target_class)
11
12      foreach class $\in$ rev_topsorted_classes do
13             class_code $\leftarrow$ getCodeForClass(class, invariants_dict)
14
15            foreach dep $\in$ getClassDependencies(class) do
16                   // Invariant: invariants have been generated for dep
17                   dep_code $\leftarrow$ getCodeForClass(dep, invariants_dict)
18                   class_code $\leftarrow$ class_code + dep_code
19
20
22
23
24      return invariants_dict[target_class.id]
25
26
27Function getCodeForClass(class, invariants_dict):
28       class_text $\leftarrow$ class.get_declaration_text()
29        // Header
31             // Get generated invariants
32             class_text $\leftarrow$ invariants_dict[class.id] + class_text
33
34      return class_text
35
Algorithm 1 Function to generate invariants for target class AST
3.1.2 Generation by LLM
After building the source program AST, ClassInvGen uses LLMs to analyze the class module and infers both invariants for the target class and tests that exercise the class’s implementation as thoroughly as possible. ClassInvGen uses a fixed system prompt that defines class invariants and outlines two main tasks: (1) generating class invariants from the source code, and (2) creating a test suite of valid API calls without specifying expected outputs (see prompt details in the Appendix 0.A). Next, ClassInvGen instantiates a user prompt template with the actual target class.
From the source program AST, ClassInvGen identifies program dependencies and populates the prompt template with the leaf struct/class. Starting from the leaf nodes, ClassInvGen leverages previously generated invariants by including them in the prompt for later classes. To accommodate the LLM’s context window limit, only the relevant child classes of the current target class are included in the prompt, with method implementations and private fields/methods hidden when necessary. An algorithm for this process is presented in Algorithm 1. Algorithm 1 presents the invariant generation process for a source program AST. The main function GenerateInvariant takes a target_class and leverages a caching mechanism through invariants_dict to avoid redundant computations (Line 1).The algorithm first collects dependent classes via getClassRecursively (Line 1) and sorts them using reverseToplogicalSort to ensure dependency-aware processing (Line 1). For each class in the sorted order, it constructs the necessary context by obtaining the class code through getCodeForClass (Line 1). For each dependency dep of the current class, it retrieves the dep_code and concatenates it with class_code (Lines 11).
The algorithm then generates invariants using generateInvariantWithLLM and stores them in invariants_dict (Line 1).
The helper function getCodeForClass constructs class representations by combining the declaration text with any existing invariants from invariants_dict (Lines 11). The algorithm concludes by returning final_invariants for the target class, effectively managing the invariant generation process while respecting LLM context limitations.
For the AvlTree example, we begin by instantiating the prompt with Node for annotation, followed by AvlTree, since Node is a subclass of AvlTree, as illustrated in Figure 3. In this specific case, however, Algorithm 1 does not make a difference due to the small size of the source program; the entire AvlTree code fits within the LLM’s context window easily.
@@ 20 T data;
Figure 3: Header file of AvlTree.
@@
In contrast, when working with the bdd_manager class in Z3 (around 1700 lines of code), ClassInvGen begins generation with bdd, a subclass of bdd_manager. Algorithm 1 enables ClassInvGen to partition the bdd_manager class and outputs meaningful class invariants (see Section 5).
3.1.3 Instrumentation
To check candidate invariants, each public method is automatically instrumented with a check_invariant call at both the start and end of its implementation. This allows us to verify that invariants hold both before and after method execution. Each invariant is implemented as a method call to prevent conflicts with local variables.
When a specific invariant is being checked, its code is plugged into the check_invariant function with assertions. This ensures that during pruning, whenever a public API call is made, each invariant candidate is automatically verified.
Additional examples of instrumentation and invariant checking are provided in Appendix 0.C.
3.2 Heuristic Pruning
@@
The LLM generates test suites that serve as filters for invariant candidates. To select the most effective test suite, we use line coverage as a metric, as it provides a straightforward proxy for test suite completeness. The test suite with the highest coverage becomes our set of filtering tests.
When generating tests, ClassInvGen creates valid sequences of API calls without asserting expected behavior, since our goal is to filter invariant candidates rather than test the source program directly. Among all generated test suites, we compile and run each one with the source program, selecting the one with the highest line coverage as the filtering tests.
@@
If an invariant candidate successfully compiles and runs with the filtering tests, it is designated a filtered invariant and included in the final output of ClassInvGen.
3.3 Refinement
For invariants that fail during compilation or runtime, ClassInvGen implements a feedback-driven refinement process. For each failing invariant, the system collects compiler output, error messages, and test results, and then feeds this information back to the LLM using a dedicated prompt template.
This feedback loop allows ClassInvGen to repair failing invariants by providing the LLM with specific error information and the context in which the error occurred. We set a default threshold of 3 refinement attempts per invariant, balancing the cost of LLM calls with the benefit of repairs.
Refinement allows ClassInvGen to fix common issues such as type errors, undefined references, and logical inconsistencies. More detailed examples of the refinement process are provided in Appendix 0.C.
4 Results
4.1 Benchmark
Table 1: Characteristics of the benchmark data structures. # LoC represents lines of code, # methods indicates the number of implemented methods, and # dep. shows the number of dependent classes for each data structure.